# LlamaFactory: Unified Efficient Fine-Tuning of 100+ Language Models

### v2

@@ 4 Empirical Study
Gemma-2B Llama2-7B Llama2-13B
Method Memory Throughput PPL Memory Throughput PPL Memory Throughput PPL
(GB) (Tokens/s) (GB) (Tokens/s) (GB) (Tokens/s)
Baseline / / 11.83 / / 7.53 / / 6.66
Full-tuning 17.06 3090.42 10.34 38.72 1334.72 5.56 / / /
Freeze-tuning 8.10 5608.49 11.33 15.69 2904.98 6.59 29.02 1841.46 6.56
GaLore 10.16 2483.05 10.38 15.43 1583.77 5.88 28.91 956.39 5.72
LoRA 7.91 3521.05 10.19 16.32 1954.07 5.81 30.09 1468.19 5.75
QLoRA 5.21 3158.59 10.46 7.52 1579.16 5.91 12.61 973.53 5.81
@@ 4.1 Training Efficiency
We utilize the PubMed Canese and Weis (2013) dataset, which comprises over 36 million records of biomedical literature. We extract around 400,000 tokens from the abstract of the literature to construct the training examples. We fine-tune the Gemma-2B (Mesnard et al., 2024), Llama2-7B and Llama2-13B (Touvron et al., 2023b) models using the generative pre-training objective with various efficient fine-tuning methods. We compare the results of full-tuning, freeze-tuning, GaLore, LoRA and 4-bit QLoRA. After fine-tuning, we calculate the perplexity on the training examples to evaluate the efficiency of different methods. We also incorporate the perplexities of the pre-trained models as baselines. More experimental details can be found in Appendix D.1.
@@ Results
The training efficiency results are presented in Table 3, where memory refers to the peak memory consumed during training, throughput is calculated as the number of tokens trained per second, and PPL represents the perplexity of the model on the training examples. Since full-tuning Llama2-13B lead to a memory overflow, the results are not recorded. We observe that QLoRA consistently has the lowest memory footprint because the pre-trained weights are represented in lower precision. LoRA exhibits higher throughput leveraging the optimization in LoRA layers by Unsloth. GaLore achieves lower PPL on large models while LoRA advantages on smaller ones.
4.2 Fine-Tuning on Downstream Tasks
Experimental Setup
To evaluate the effectiveness of different efficient fine-tuning methods, we compare the performance of various models after fine-tuning on downstream tasks. We construct the training set and test set using 2,000 examples and 1,000 examples from three representative text generation tasks, including CNN/DM (Nallapati et al., 2016), XSum (Narayan et al., 2018) and AdGen (Shao et al., 2019), respectively. We select several instruction-tuned models and fine-tune them following the sequence-to-sequence task using different fine-tuning methods. We compare the results of full-tuning (FT), GaLore, LoRA and 4-bit QLoRA. After fine-tuning, we calculate the ROUGE score (Lin, 2004) on the test set of each task. We also incorporate the scores of the original instruction-tuned models as baselines. More experimental details can be found in Appendix D.2.
Results
The evaluation results on downstream tasks are shown in Table 4. We report the averaged scores over ROUGE-1, ROUGE-2 and ROUGE-L for each LLM and each dataset. Some results of Gemma-7B are not included in the table because the GaLore method is not applicable to this model. An interesting finding from the results is that LoRA and QLoRA achieve the best performance in most cases, except for the Llama2-7B and ChatGLM3-6B models on the CNN/DM and AdGen datasets. This phenomenon highlights the effectiveness of these efficient fine-tuning methods in adapting LLM models to specific tasks. Additionally, we observe that the Mistral-7B model performs better on English datasets while the Qwen1.5-7B model achieves higher scores on Chinese dataset. These results suggest that the performance of the fine-tuned models is also associated with their inherent capabilities on specific languages.
@@ 5 Conclusion and Future Work
In the future, we will consistently keep LlamaFactory synchronous with the state-of-the-art models and efficient fine-tuning techniques. We also welcome contributions from the open-source community. In future versions, we will explore more advanced parallel training strategies and multimodal efficient fine-tuning of LLMs.
@@ 6 Broader Impact and Responsible Use
LlamaFactory has attracted a large number of individuals interested in LLMs to explore the possibility of fine-tuning their own models. This contributes significantly to the growth of the open-source community. It is gaining increasing attention and is being featured in Awesome Transformers [^3] https://github.com/huggingface/transformers/blob/v4.39.0/awesome-transformers.md as a representative of efficient fine-tuning frameworks for LLMs. We anticipate that practitioners build their LLMs upon our framework that bring benefits to society. Adherence to the model license is mandatory when using LlamaFactory for fine-tuning LLMs, thus preventing from any potential misuse.
Appendix A Related Work
In this section, we enumerate existing frameworks for fine-tuning LLMs, especially highlight those for efficient fine-tuning.
LLaMA-Adapter (Zhang et al., 2022) efficiently fine-tunes the Llama model (Touvron et al., 2023a) for instruction-following abilities using zero-init attention. FastChat (Zheng et al., 2023) is a framework focused on training and evaluating LLMs for chat completion purposes. LitGPT (AI, 2023) provides the implementation of generative models and supports various training methods. Open-Instruct (Wang et al., 2023d) is a fine-tuning framework that supports efficient fine-tuning. Colossal AI (Li et al., 2023b) takes advanced parallelism strategies for distributed training. LMFlow (Diao et al., 2023) is an extensible and efficient fine-tuning framework that supports decoder-only models, supporting both full-tuning and adapter tuning. GPT4All (Anand et al., 2023) allows LLMs to run on consumer devices, while also providing fine-tuning capabilities.
In contrast to the frameworks presented above, LlamaFactory supports a broader range of efficient fine-tuning methods. We compare our framework with existing work in Table 5.
Table 6: List of supported models.
Model Variant Organization
Llama (Touvron et al., 2023a) 7B/13B/33B/65B Meta AI
Llama 2 (Touvron et al., 2023b) 7B/13B/70B Meta AI
Baichuan Yang et al. (2023) 7B/13B Baichuan Inc
Baichuan2 (Yang et al., 2023) 7B/13B Baichuan Inc
BLOOM (Le Scao et al., 2022) 560M/3B/7.1B BigScience
BLOOMZ (Le Scao et al., 2022) 560M/3B/7.1B BigScience
ChatGLM2 (Du et al., 2022) 6B THUDM
ChatGLM3 (Du et al., 2022) 6B THUDM
ChineseLLaMA2 (Cui et al., 2023) 3B/7B/13B HFL
DeepSeek-LLM (Bi et al., 2024) 7B/67B DeepSeek
DeepSeek-Math (Shao et al., 2024) 7B DeepSeek
DeepSeek-MoE (Dai et al., 2024) 16B DeepSeek
DeepSeek-Coder (Guo et al., 2024) 6.7B/7B/33B DeepSeek
Falcon (Almazrouei et al., 2023) 7B/40B/180B TII
Gemma (Mesnard et al., 2024) 2B/7B Google
InternLM (Team, 2023) 7B/20B Shanghai AI Lab
InternLM2 (Team, 2023) 1.8B/7B/20B Shanghai AI Lab
Mistral (Jiang et al., 2023a) 7B Mistral AI
Mixtral (Jiang et al., 2023a) 8x7B Mistral AI
OLMo (Groeneveld et al., 2024) 1B/7B Allen AI
OpenChat (Wang et al., 2023b) 7B OpenChat
Orion (Chen et al., 2024a) 14B OrionStar
Phi-1.5 (Li et al., 2023c) 1.3B Microsoft
Phi-2 (Li et al., 2023c) 2.7B Microsoft
Qwen (Bai et al., 2023) 1.8B/7B/14B/72B Alibaba Cloud
Qwen1.5 (Bai et al., 2023) 1.8B/7B/14B/72B Alibaba Cloud
SOLAR (Kim et al., 2023) 10.7B Upstage AI
Skywork (Wei et al., 2023) 13B Skywork
StarCoder2 (Lozhkov et al., 2024) 3B/7B/15B BigCode
Vicuna1.5 (Zheng et al., 2023) 7B/13B LMSYS
Yi (Young et al., 2024) 6B/9B/34B 01.AI
Yuan2 (Wu et al., 2023) 2B/51B/102B IEIT
Zephyr (Tunstall et al., 2023) 7B Hugging Face H4
Appendix B Supported Models in LlamaFactory
We select the popular pre-trained models supported by LlamaFactory and list them in Table 6, ranging from dense models to sparse mixture-of-expert (MoE) models.
Table 7: Prompt of function calling in LlamaFactory.
You have access to the following tools:
> Tool Name: tool1
Tool Description: the usage of tool1
Tool Args:
- arg1 (dtype): the usage of arg1
Use the following format if using a tool:
Action: tool name (one of [tool1]).
Action Input: the input to the tool, in a JSON format representing the kwargs (e.g. ```{"input": "hello world", "num_beams": 5}```).
Appendix C Details of Designing Chat Template
Although Transformers 4.34.0 introduces the feature of chat templates, several models (such as ChatGLM and Qwen) have tokenizers that exhibit specific behaviours and cannot directly encode the special tokens. Therefore, we design a Formatter class to robustly convert the textual inputs to their embedding IDs. Specifically, we provide the EmptyFormatter, StringFormatter, FunctionFormatter and ToolFormatter. Furthermore, LlamaFactory supports fine-tuning models to obtain function calling capabilities. While the ReAct prompting Yao et al. (2023) is a popular choice for tool using, it is insufficient for nested tool parameters. We optimize the tool prompts and provided them in Table 7.
Appendix D Experiment Details
D.1 Training Efficiency
We adopt a learning rate of $10^{-5}$, a token batch size of $512$ to fine-tune these models using the 8-bit AdamW optimizer (Dettmers et al., 2022b) in bfloat16 precision with activation checkpointing to reduce the memory footprint. In freeze-tuning, we only fine-tune the last $3$ decoder layers of the model. For GaLore, we set the rank and scale to $128$ and $2.0$, respectively. For LoRA and QLoRA, we attach adapters to all linear layers and set the rank and alpha to $128$ and $256$, respectively. All the experiments are conducted on a single NVIDIA A100 40GB GPU. We enable flash attention in all experiments and Unsloth for LoRA and QLoRA experiments.
D.2 Fine-Tuning on Downstream Tasks
We briefly describe the datasets used for evaluating the effectiveness of different fine-tuning methods.
  • CNN/DM (Nallapati et al., 2016) This dataset is used for the summary generation task. English articles are gathered from CNN and Daily Mail. We utilize version 1.0.0, which comprises over 300k samples.
  • XSum (Narayan et al., 2018) This dataset is used for the abstractive summary generation task, which consists of English documents and one-sentence summaries, totaling over 200k samples.
  • AdGen (Shao et al., 2019) This dataset is a Chinese dataset that requires models to generate advertising text based on given keywords, totalling over 100k samples.
In this experiment, we adopt a learning rate of $10^{-5}$. We set batch size to $4$ while the maximum input length is $2048$. We fine-tune the models using the 8-bit AdamW optimizer (Dettmers et al., 2022b) in bfloat16 precision with activation checkpointing to reduce the memory footprint. For GaLore, we set the rank and scale to $128$ and $2.0$, respectively. For LoRA and QLoRA, we attach adapters to all linear layers and set the rank and alpha to $128$ and $256$, respectively. All the experimental results are carried out on the NVIDIA A100 40GB GPUs with flash attention enabled.

### v3

@@ 5 Empirical Study
5.1 Training Efficiency
Experimental Setup
We utilize the PubMed dataset Canese and Weis (2013), which comprises over 36 million records of biomedical literature. We extract around 400K tokens from the abstract of the literature to construct the training corpus. Then we fine-tune the Gemma-2B (Mesnard et al., 2024), Llama2-7B and Llama2-13B (Touvron et al., 2023b) models using the generative pre-training objective with various efficient fine-tuning methods. We compare the results of full-tuning, freeze-tuning, GaLore, LoRA and 4-bit QLoRA. After fine-tuning, we calculate the perplexity on the training corpus to evaluate the efficiency of different methods. We also incorporate the perplexities of the pre-trained models as baselines.
In this experiment, we adopt a learning rate of $10^{-5}$, a token batch size of $512$. We fine-tune these models using the 8-bit AdamW optimizer (Dettmers et al., 2022b) in bfloat16 precision with activation checkpointing to reduce the memory footprint. In freeze-tuning, we only fine-tune the last $3$ decoder layers of the model. For GaLore, we set the rank and scale to $128$ and $2.0$, respectively. For LoRA and QLoRA, we attach adapters to all linear layers and set the rank and alpha to $128$ and $256$, respectively. All the experiments are conducted on a single NVIDIA A100 40GB GPU. We enable flash attention in all experiments and Unsloth for LoRA and QLoRA experiments.
Results
The results about the training efficiency are presented in Table 4, where memory refers to the peak memory consumed during training, throughput is calculated as the number of tokens trained per second, and PPL represents the perplexity of the model on the training corpus. Since full-tuning Llama2-13B lead to a memory overflow, the results are not recorded. We observe that QLoRA consistently has the lowest memory footprint because the pre-trained weights are represented in lower precision. LoRA exhibits higher throughput leveraging the optimization in LoRA layers by Unsloth. GaLore achieves lower PPL on large models while LoRA advantages on smaller ones.
5.2 Fine-Tuning on Downstream Tasks
@@
To evaluate the effectiveness of different efficient fine-tuning methods, we compare the performance of various models after fine-tuning on downstream tasks. We construct non-overlapping training set and test set using 2,000 examples and 1,000 examples from three representative text generation tasks, including CNN/DM (Nallapati et al., 2016), XSum (Narayan et al., 2018) and AdGen (Shao et al., 2019), respectively. We select several instruction-tuned models and fine-tune them following the sequence-to-sequence task using different fine-tuning methods. Then we compare the results of full-tuning (FT), GaLore, LoRA and 4-bit QLoRA. After fine-tuning, we calculate the ROUGE score (Lin, 2004) on the test set of each task. We also incorporate the scores of the original instruction-tuned models as baselines.
In this experiment, we set learning rate to $10^{-5}$, batch size to $4$ and maximum input length to $2048$. We fine-tune these models using the 8-bit AdamW optimizer (Dettmers et al., 2022b) in bfloat16 precision with activation checkpointing. For GaLore, we set the rank and scale to $128$ and $2.0$, respectively. For LoRA and QLoRA, we attach adapters to all linear layers and set the rank and alpha to $128$ and $256$, respectively. All the experiments are conducted on NVIDIA A100 40GB GPUs.
@@ Results
The evaluation results on downstream tasks are shown in Table 5. We report the averaged scores over ROUGE-1, ROUGE-2 and ROUGE-L. Some results of the Gemma-7B and Qwen2-7B (Bai et al., 2023) models are not included in the table because the GaLore method may not be applicable to them. An interesting finding from the results is that LoRA and QLoRA achieve the best performance in most cases, except for the ChatGLM3-6B (Zeng et al., 2024) and Llama2-7B models on the CNN/DM and AdGen datasets. This phenomenon highlights the effectiveness of these efficient fine-tuning methods in adapting LLMs to specific tasks. Additionally, we observe that Llama3-8B achieves the best performance among these models, while Yi-6B (Young et al., 2024) and Mistral-7B (Jiang et al., 2023a) exhibit competitive performance among models of the same size.
@@ 6 Conclusion and Future Work
We will consistently keep LlamaFactory synchronous with the state-of-the-art models and efficient fine-tuning techniques. We also welcome contributions from the open-source community. The road map of LlamaFactory including:
(1) Enabling fine-tuning for models that supports a wider range of modalities, e.g., the audio and video modalities (Zhu et al., 2024a).
(2) Integrating more parallel training strategies, e.g., sequence parallelism (Jacobs et al., 2023) and tensor parallelism (Shoeybi et al., 2019).
(3) Exploring stronger fine-tuning methods for conversational models, e.g., self-play (Chen et al., 2024c; Yuan et al., 2024).
@@ 7 Broader Impact and Responsible Use
LlamaFactory has attracted a large number of individuals interested in LLMs to explore the possibility of customizing models. This contributes significantly to the growth of the open-source communities. It is gaining increasing attention and is being featured in Awesome Transformers [^2] https://github.com/huggingface/transformers/blob/v4.40.0/awesome-transformers.md#llama-factory as a representative of efficient fine-tuning frameworks for LLMs. We anticipate that practitioners build their LLMs upon our framework that bring benefits to society. Adherence to the model license is mandatory when using LlamaFactory for fine-tuning LLMs, thus preventing from any potential misuse.
Acknowledgements
This work is supported partly by the National Science and Technology Major Project under Grant 2022ZD0120202, by the National Natural Science Foundation of China (No. U23B2056), by the Fundamental Research Funds for the Central Universities, and by the State Key Laboratory of Complex & Critical Software Environment.